HUBBLE SPACE TELESCOPE DAILY REPORT #1684 PERIOD COVERED: 0000Z (UTC) 07/26/96 - 0000Z (UTC) 07/29/96 Daily Status Report as of 211/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/1/2 6665 (Galaxy Rotation Curves at Large Radius = Using Lyman-Alpha Absorption Line) The High Resolution Spectrograph (Both Sides One and Two) was used = to make spectrographic observations of the background QSO PG13099+355 = through the disk of two foreground spiral galaxies in order to obtain the = galaxy's rotation curve by studying the structure of the Lyman-Alpha = lines. The observations were executed as scheduled, and no anomalies = were noted. 1.2 Completed FOS/BL 6166 (FOS Cycle 5: Location of Spectra and = Wavelength Calibration) The Faint Object Spectrograph (Blue Side) was used to make = wavelength calibration images to measure the stability of the FOS Blue = Side wavelength scale and spectral locations. A Taled calibration = observation was also taken. The observations were completed as planned, = and no anomalies were reported. 1.3 Completed FOS/BL 6663 (Systematic Investigation of C/1995 01 = (Hale-Bopp)) The Faint Object Spectrograph (Blue Side) was used to make = observations of the comet Hale-Bopp (aka, comet C/1995-01) as part of a = systematic investigation of the comet. The comet is being monitored over = a wide range of heliocentric distances. These data at different = distances will provide unique information the sublimation behavior of the = material in this comet. The observations were completed as planned, and = no problems were reported. 1.4 Completed Two Sets of WF/PC-2 6903 (WF/PC-2 Cycle 6 Calibration: = Decontamination) The WF/PC-2 was used to make several calibration bias = observations, internal flats, K-Spot images and dark calibration images = which are all part of a Cycle 6 decontamination procedure. No science = was interleaved with the dark exposures. The observations were completed = as scheduled, and no problems were reported. 1.5 Completed Two Sets of FOS/RD 6204 (FOS Cycle 5 Internal/External = Offset) The Faint Object Spectrograph (Red Side) was used to make = spectrographic calibration images of the objects NGC-6833 and HD207757 in = order to measure the spectrograph's internal and external offsets. The = internal lamp was intermingled with the external observations. During = the second iteration of this proposal, between approximately 210/1435Z = and 210/1510Z, there were repeated losses of FGS lock which, in turn, = caused an FOS target acquisition failure. These events are documented in = HSTAR 1996. Otherwise, the observations were completed as planned, and = no further problems were reported. 1.6 Completed Two Sets of FOC/96 6349 (Snapshot search for High Redshif= t QSOs with Far-UV Flux) The Faint Object Camera (f/96) was used to make snapshot images of = the quasi-stellar objects Q1158-187 and Q0109-353. This is to determine = if these high redshifted objects have a far ultraviolet flux. The = observations were executed as scheduled, and no anomalies were noted. 1.7 Completed FGS 6885 (L726-8/ UV-Ceti) The FGSs were used in both the POS and TRANS modes to make = astrometric observations of the star UV-Ceti. This is to look for short = period, low mass companions in orbit around the star. The observations = were completed as planned, an no anomalies were reported. 1.8 Completed FGS 6270 (The Parallax of RW Tri) The FGS were used in the POS Mode to make astrometric observations = of old nova RW Tri and several other reference stars to measure its = parallax. The observations were completed as planned, and no anomalies = were noted. 1.9 Completed FOS/BL 6476 (A Study of Curcumstellar Material Formed in = a Hydrogen-Deficient Environment) The Faint Object Spectrograph (Blue Side) was used to make low = resolution ultraviolet spectral observations of the hot variable CrB-type = star V348-SGR in order to study the curcumstellar material formed in its = hydrogern-defecient environment. The objective is to study the newly = formed dust associated with the brightness changes. The observations were = completed as planned, and no problems were reported. 1.10 Completed FGS 6049 (Monitoring the Size and Shape of Mira Type = Stars) The FGSs were used to make astrometric observations of the Mira = class star HD120285 in order to study its size and shape during different = phases of its pulsation cycle. The observations were completed as = planned, and no anomalies were reported. 1.11 Completed FOS/BL 6470 (Decoding the Inner Disk About = Beta-Pictoris) The Faint Object Spectrograph (Blue Side) was used to make = ultraviolet spectrographic observations of the disk around the star Beta = Pictoris (aka, HD50241) looking for emission lines arising in the inner = disk. The objective is to ascertain the composition of the inner disk. = The observations were completed as planned, and no problems were = reported. 1.12 Completed FOS/RD 6925 (FOS Observations of Solar Analog = Standards) The Faint Object Spectrograph (Red Side) was used to make absolute = spectrophotmetry observations of the solar analog star P303E-Cali in the = spectral range of 2200 A0 to 8500 A0. The objective is to provide the = primary standards for the absolute calibration of the NICMOS. The = observations were completed as scheduled, and no problems were reported. 1.13 Completed WF/PC-2 6717 (UV Spectroscopy of the Luminous Cooling = Flow Nebula in A2597) The WF/PC-2 was used to make ultraviolet observations of the = nuclear regions within the luminous, cooling flow, nebula A2597. The = observations were completed as planned and no problems were reported. 1.14 Completed FOS/RD 6548 (The Cool Interstellar Medium in Hot = Cluster of Galaxies) The Faint Object Spectrograph (Red Side) was used to make = ultraviolet resonance absorption line observations of quasar QSO0107-15 = in order to study the cool interstellar medium within a cluster of = galaxies which lie within the line of sight to the quasar. The QSO is = being used a probe of the galactic cluster. The observations were = executed as scheduled, and no anomalies were noted. 1.15 Completed WF/PC-2 6902 (WF/PC-2 Cycle 6 Photometric Calibration = Monitor UV/OPT STD) The WF/PC-2 was used to make both ultraviolet and optical = photometric observations of the standard star GRW+70D582 as part of a = Cycle 6 calibration procedure. This observation was done immediately = before and after decontamination procedures. The observations were = completed as planned, and no problems were reported. 1.16 Completed FOS/RD 6595 (Does Pluto Have a Haze Layer?) The Faint Object Spectrograph (Red Side) was used to make = spectrographic observations of the planet Pluto in order to determine if = the planet has a haze layer. These observations were made during a = series of stellar occultations. The objective is to study the physical = structure of the atmosphere at kilometer-level spatial resolution. The = observations were executed as scheduled, and no anomalies were noted. 1.17 Completed FOS/RD 6585 (UV Spectra of Representative Stellar = Populations of Age > 2) The Faint Object Spectrograph (Red Side) was used to make = mid-ultraviolet spectrographic observations of nuclear region of nearby = galaxy NGC-5831. This galaxy has a relatively old stellar population. = The objective is to fill in crucial gaps in our understanding of stellar = populations, and thus aid in the modeling of galactic evolution. The = observations were completed as planned, and no anomalies were reported. 1.18 Completed FGS 6641 (Continuation of Mass Ratio Determination for = Double Stars) The Fine Guidance Sensors were used in both the POS and TRANS = modes to make astrometric observations of the double star L722-22 in = order to determine the mass ratio of its two components. This binary = system contains either a low-mass star or a brown dwarf. The = observations were executed as scheduled, and no anomalies were noted. 1.19 Completed WF/PC-2 6346 (Snapshot Survey of the Ultraluminous IRAS = Galaxy Sample-Part 1 High) The WF/PC-2 was used to make snapshot survey observations of the = ultraluminous IRAS galaxy CL-ULIR234. These data will be searched for = multiple cores, which imply a merger has occurred, and for starburst = regions. The observations were executed as scheduled, and no anomalies = were noted. 1.20 Completed WF/PC-2 6348 (Blue Continuum Snapshots of 3CR Radio = Galaxies) The WF/PC-2 was used to make some snapshot images of the blue = continuum in radio galaxy 3C2.0. Combined with existing data, these = exposures will provide high spatial resolution information on the color = distribution within galaxies, redding due to dust, spacial indices of the = optical synchrotron jets, and colors of the active nuclei and tidal jets. = The observations were executed as scheduled, and no anomalies were = noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 22 Successful : 22 As more completely described in 1.5, there were numerous losses of = lock during the interval 210/1435-1510Z. Scheduled Reacquisitions : 22 Successful : 22 2.2 FHST Updates: Scheduled : 39 Successful : 39 2.3 Operations Notes: All science data was collected from the previous week's = observations. Most TDRSS services were manually scheduled. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH